Question 2 1/3 pts Let S be the event that a randomly selected college student has taken a statistics course, and C be the event that the student has taken a chemistry course. Suppose P(S) 0.4, P(C) -0.3 and P(Snc) 0.2 Suppose a student is randomly selected. (a) Find the probability that the student has taken Statistics, Chemistry, or both. I Select ] (b) Find the probability that the student has taken neither statistics nor chemistry. I Select ] (c) Find the probability that the student has taken Statistics, but not Chemistry I Select ]
Answer #1

(a)

Desired Probability is,

P(SU C) P(S) P(C) _ P(S C) 0.4 + 0.3-0.2 0.5 (Ans)

(b)

Desired Probability is,

P(S U C)c-1-P(SU C)-1-0.5 = 0.5 (Ans)

(c)

Desired Probability is,

P(S) _ P(S C) 0.4-0.2 = 0.2 (Ans)

** If the answer does not match then please comment.